A girl buys three apples and seven oranges for $4.56. If an orange cost .28 more than an apple, how much does each fruit cost?
Answer by checkley77(12844)

3X+7(X+.28)=4.56
3X+7X+1.96=4.56
10X=4.56-1.96
10X=2.60
X=2.60/10
X=.26 COST OF AN APPLE.
.26+.28=.54 COST OF AN ORANGE.
PROOF:
3*.26+7*.54=4.56
.78+3.78=4.56
4.56=4.56